# Preferential Attachment Trees with Vertex Death: Persistence of the Maximum Degree

### v1

@@ 5.2. The oldest alive individual
In this sub-section we study the quantity $O_{t}^{\mathrm{ct}}$; the birth-time of the oldest alive individual in $\mathrm{BP}(t)$ conditionally on survival. To this end, we first obtain precise asymptotic distributional properties of the lifetime $L=S_{D+1}$ of individuals in the branching process, to then prove when the oldest individual alive at time $t$ is born.
@@ 5.2.1. Lifetime distribution
We know from [23, Lemma $6.1$] that the lifetime $L$ of an individual is asymptotically distributed as an exponential random variable with rate $d^{*}$ when $d$ converges to $d^{*}\leq R$. That is $\mathbb{P}\!\left(L>t\right)=\exp(-(d^{*}+o(1))t)$ as $t$ tends to infinity. The following result provides a higher-order result for the tail distribution when Assumption (C-V) is also satisfied.
@@
Remark 5.7.
This result complements [23, Lemma $6.2$] where a different third-order rescaling is provided for $\log\mathbb{P}\!\left(L>t\right)$ in the case that Assumption (C-V) is not satisfied.
@@
We start by making the following observations. Since $\varphi_{2}$ converges and $d$ converges, it follows that $\rho_{2}$, as in (5.1), converges. Furthermore, as $\varphi_{2}(k)={\rm Var}(S_{k})$ converges and $\varphi_{1}(k)=\mathbb{E}\left[S_{k}\right]$, it follows that $S_{k}-\varphi_{1}(k)$ converges almost surely to some random variable $M_{\infty}$.
Lower bound. We set $k:=\lfloor\varphi_{1}^{-1}(t)\rfloor$. Then, using Lemma 5.1 and using the above observations,
@@
$=\mathrm{e}^{-\rho_{1}(k)+\mathcal{O}(1)}\mathbb{P}\!\left(S_{k}>t\right)$
$\geq\mathrm{e}^{-\rho_{1}(\varphi_{1}^{-1}(t))+\mathcal{O}(1)}\mathbb{P}\!\left(S_{k}-\varphi_{1}(k)>0)\right)$
$=\mathrm{e}^{-d^{*}t-\mathcal{K}_{\alpha}(t)+\mathcal{O}(1)}(\mathbb{P}\!\left(M_{\infty}>0\right)+o(1)),$
@@
$M:=\lceil(1-\varepsilon)t\rceil\qquad\text{and}\qquad k_{i}:=\lceil\varphi_{1}^{-1}(t-i)\rceil\quad\text{for }i\in\{0,\ldots,M\}.$ (5.15)
@@
As $\theta(1-\varepsilon)>d^{*}$ by the choice of $\theta$ and $\varepsilon$, and $\mathcal{K}_{\alpha}=o(t)$ by Lemma 5.5 $(c)$, it follows that the second term is negligible compared to the first term. To deal with the sum, we define
$\widetilde{\alpha}(k)=\sum_{i=0}^{k-1}\frac{|d(i)-d^{*}|}{b(i)+d(i)}\qquad\text{for }k\in\mathbb{N}_{0},\quad\text{and}\quad\mathcal{K}_{\widetilde{\alpha}}(t):=\widetilde{\alpha}\big{(}\varphi_{1}^{-1}(t)\big{)}\qquad\text{for }t\geq 0,$ (5.19)
@@
We note that this is possible since $\mathcal{K}_{\widetilde{\alpha}}(t)=o(t)$ (this follows in the same way as $\mathcal{K}_{\alpha}(t)=o(t)$ in Lemma 5.5 $(c)$) and $M\geq(1-\varepsilon)t$. Using the second part of Lemma 5.5 $(c)$ yields
@@
$-\mathcal{K}_{\alpha}(t-(i+1))\leq\sup_{x\in[\varepsilon t,t]}|\mathcal{K}_{\alpha}(x)|\leq\sup_{x\in[\varepsilon t,t]}\mathcal{K}_{\widetilde{\alpha}}(x)=\mathcal{K}_{\widetilde{\alpha}}(t),$ (5.24)
@@ 5.2.2. Non-survival of old individuals
$\ell_{t}:=\frac{d^{*}}{\lambda^{*}+d^{*}}t+\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))-K,\qquad u_{t}:=\frac{d^{*}}{\lambda^{*}+d^{*}}t+\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))+K.$ (5.28)
@@
$\frac{M}{\mathbb{P}\!\left(\mathcal{S}\right)}\exp\big{(}\lambda^{*}{}$ $\ell_{t}-d^{*}(t-A-\ell_{t})-\mathcal{K}_{\alpha}(t-A-\ell_{t})+C_{1}\big{)}$ (5.31)
$=\frac{M}{\mathbb{P}\!\left(\mathcal{S}\right)}\exp\Big{(}{}$ $\mathcal{K}_{\alpha}(r(t))-\mathcal{K}_{\alpha}\Big{(}\frac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))+K-A\Big{)}-(\lambda^{*}+d^{*})K+d^{*}A+C_{1}\Big{)},$
where $C_{1}>0$ is a sufficiently large constant. We use Lemma 5.5 $(c)$ to obtain that for any $K,A>0$ there exists $t^{\prime}>0$ such that
@@
That is, the left-hand side is at most $1$ for all $t\geq t^{\prime}=t^{\prime}(K,A)$. Then, Assumption $\mathcal{K}_{\alpha}$ yields that there exists a constant $C_{2}>0$ such that for all $t>0$,
@@
$\mathbb{P}_{\mathcal{S}}\left(O^{\mathrm{ct}}_{t-A}\leq\ell_{t}\right)\leq\frac{M}{\mathbb{P}\!\left(\mathcal{S}\right)}\exp\big{(}C_{2}+1-(\lambda^{*}+d^{*})K+d^{*}A+C_{1}\big{)}+\varepsilon/4\leq\varepsilon/2,$ (5.34)
where we choose $K$ sufficiently large so that the first term is at most $\varepsilon/4$ in the final step. We note that we can first choose such a $K$ and then take $t^{\prime}=t^{\prime}(K,A)$ sufficiently large such that (5.32) holds.
@@
$\mathcal{E}_{t}:=\{\forall v\in\mathcal{B}(\widetilde{u}_{t},u_{t}):L^{(v)}\leq t+A-\widetilde{u}_{t}\},\qquad\mathcal{D}_{t}:=\{|\mathcal{B}(\widetilde{u}_{t},u_{t})|>m\}\quad\text{for }m\in\mathbb{N},t\geq 0.$ (5.36)
@@
$\mathbbm{1}_{\mathcal{D}_{t}}\mathbb{P}\!\left(\mathcal{E}_{t}\,|\,\mathcal{F}_{\widetilde{u}_{t}},|\mathcal{B}(\widetilde{u}_{t},u_{t})|\right)\leq\mathbbm{1}_{\mathcal{D}_{t}}\big{(}1-\mathbb{P}\!\left(L\geq\overline{r}(t)-\widetilde{u}_{t}\right)\big{)}^{m}\leq\exp\big{(}-m\mathbb{P}\!\left(L\geq\overline{r}(t)-\widetilde{u}_{t}\right)\big{)},$ (5.39)
@@
We choose $m=M^{-1}\exp(\lambda^{*}u_{t})$ for a large constant $M>0$ to be determined to obtain the upper bound
@@
where $C_{3}>0$ is a large constant. We again use Lemma 5.5 $(c)$ and Assumption $\mathcal{K}_{\alpha}$ to obtain similar bounds as in (5.32) and (5.33). We thus obtain, for some constants $C_{3}>0$ and $t^{\prime}=t^{\prime}(K,A)>0$ that for all $t>t^{\prime}$ we have have the upper bound
$\mathbb{P}(\mathcal{S})^{-1}\exp\Big{(}-M^{-1}\exp\Big{(}C_{3}+(d^{*}/2+\lambda^{*})K-d^{*}A-C_{3}\Big{)}\Big{)}$ (5.43)
bound the first two terms in the exponential from below by $-C_{4}$, uniformly in $A$ and $K$ and for all $t>t^{\prime}$, where $t^{\prime}>0$ and $C_{4}>0$ are large constants. Then, we choose $K$ sufficiently large so that the entire term is at most $\varepsilon/4$. Again, we can first choose such a $K$ and then choose $t^{\prime}$.
@@
[…]
@@ 5.3. The richest individual
In this sub-section we study $I^{\mathrm{ct}}_{t}$, the birth-time of the richest individual in $\mathrm{BP}(t)$. That is, the individual that is alive at time $t$ and has produced the largest offspring by time $t$. To this end, we show that individuals that are ‘too young’ produce ‘small’ offspring, whereas there exist ‘old’ individuals that produce large offspring.
@@ 5.3.1. Young individuals have small offspring and no individual has too large offspring
In this section we show that individuals that are alive in an interval of time centred at $t$ but born ‘too late’, that is, born too long after the oldest alive individual has been born, have too small offspring to attain the largest offspring among all alive individuals. Similarly, we show that there does not exist any individual that attains a ‘too large’ offspring by time $\approx t$.
@@
$k$ $:=\lceil\varphi_{1}^{-1}\big{(}\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\tfrac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))-C_{1}\big{)}\rceil,\qquad u_{t}:=\frac{d^{*}}{\lambda^{*}+d^{*}}t+\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t)),$ (5.48)
$k^{\prime}$ $:=\lceil\varphi_{1}^{-1}\big{(}\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\tfrac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))+C_{1}^{\prime}\big{)}\rceil.$
Then, we have the following lemma.
@@
$\mathbb{P}_{\mathcal{S}}\left(\forall s\in[t-A,t+A]\ \forall v\in\mathcal{B}(u_{t}+C_{2},t)\cap\mathcal{A}_{s}^{\mathrm{ct}}:d_{s}(v)<k\right)\geq 1-\varepsilon.$ (5.49)
@@
$\mathbb{P}_{\mathcal{S}}\left(\forall s\in[t-A,t+A]\ \forall v\in\mathcal{A}_{s}^{\mathrm{ct}}:d_{s}(v)<k^{\prime}\right)\geq 1-\varepsilon.$ (5.50)
@@
$\mathbb{P}_{\mathcal{S}}{}$ $\big{(}\exists s\in[t-A,t+A]\ \exists v\in\mathcal{B}(u_{t}+C_{2},t)\cap\mathcal{A}_{s}^{\mathrm{ct}}:d_{s}(v)\geq k\big{)}$ (5.51)
@@
We now introduce the event $\{\sup_{t\geq 0}|\mathcal{B}(0,t)|\mathrm{e}^{-\lambda^{*}t}\leq M\}$, for some large but fixed $M>0$ to be determined. We then partition the set $\mathcal{B}(u_{t},t)$ into sets of individuals born in the intervals $(u_{t}+C_{2}+(j-1),u_{t}+C_{2}+j)$, i.e. into sets $\mathcal{B}(u_{t}+C_{2}+(j-1),u_{t}+C_{2}+j)$, with $j\in[\lceil t-u_{t}-C_{2}\rceil]$ and apply a union bound. For individuals $v\in\mathcal{B}(u_{t}+C_{2}+(j-1),u_{t}+C_{2}+j)$, we can bound their birth-time $\mathcal{B}(v)$ from below and above by $u_{t}+C_{2}+(j-1)$ and $u_{t}+C_{2}+j$, respectively. Similarly we can bound $s$ from below and above by $t-A$ and $t+A$, respectively. As a result, we bound the above probability from above by
@@
We now choose $M$ large enough so that, by (5.11) in Corollary 5.4, the probability on the second line is at most $\varepsilon/2$. It thus remains to bound the sum from above by $\varepsilon/2$ to arrive at the desired result. For any $j\in[\lceil t-u_{t}\rceil]$, we omit the part of the event that $S_{D+1}>t-A-(u_{t}+C_{2}+j)$, use the independence of $D$ and $S_{k}$ and use a Chernoff bound to arrive at the upper bound
@@
$-\log\mathbb{E}\left[\mathrm{e}^{-\theta S_{k}}\right]=\sum_{i=0}^{k-1}\log\Big{(}1+\frac{\theta}{b(i)+d(i)}\Big{)}\geq\theta\varphi_{1}(k)-\frac{\theta^{2}}{2}\varphi_{2}(k).$ (5.54)
Since $\varphi_{2}(k)$ is increasing and converges to $\varphi_{2}(\infty):=\lim_{k\to\infty}\varphi_{2}(k)$ as $k$ tends to infinity by Assumption (C-V), and by using Lemma 5.1, we finally arrive at the upper bound
@@
We apply Lemma 5.5 $(c)$ and Assumption $\mathcal{K}_{\alpha}$ and set $\theta=(C_{2}+(j-1)-(A+C_{1}))/\varphi_{2}(\infty)$ (which is positive for all $j\geq 1$ when $C_{2}$ is large), to bound this for all $t>t^{\prime}$ by
@@
where $C>0$ and $t^{\prime}$ are fixed constants, where $C$ is independent of all other constants and parameters and $t^{\prime}=t^{\prime}(C_{1},d^{*},\lambda^{*})$. By substituting this upper bound in the sum in (5.52), we can bound it from above by

### v2

@@ 5.2. The oldest alive individual
In this sub-section we study the quantity $O_{t}^{\mathrm{ct}}$; the birth-time of the oldest alive individual in $\mathrm{BP}(t)$ conditionally on survival. To this end, we first obtain precise asymptotic distributional properties of the lifetime $L=S_{D+1}$ of individuals in the branching process, to then prove an asymptotic expansion for $O_{t}^{\mathrm{ct}}$ as $t\to\infty$.
@@ 5.2.1. Lifetime distribution
We know from [23, Lemma $6.1$] that the lifetime $L$ of an individual is asymptotically distributed as an exponential random variable with rate $d^{*}$ when $d$ converges to $d^{*}\leq R$. That is $\mathbb{P}\!\left(L>t\right)=\exp(-(d^{*}+o(1))t)$ as $t$ tends to infinity. A more precise asymptotic expansion is provided in [23, Lemma $6.3$] when Assumption (C-V) is not satisfied. We complement this result by the following lemma, which provides a (different) higher-order expansion when Assumption (C-V) is satisfied.
@@
We start by making the following observations. Since $\varphi_{2}$ converges and $d$ converges, it follows that $\rho_{2}$, as in (5.3), converges. Furthermore, as $\varphi_{2}(k)={\rm Var}(S_{k})={\rm Var}(S_{k}-\varphi_{1}(k))$ converges and $\varphi_{1}(k)=\mathbb{E}\left[S_{k}\right]$, it follows that $S_{k}-\varphi_{1}(k)$ converges almost surely to some random variable $M_{\infty}$ by Kolmogorov’s two series theorem [19, Theorem 2.5.6].
Lower bound. Set $k\coloneq\lfloor\varphi_{1}^{-1}(t)\rfloor$. Then, using Lemma 5.1 and the above observations,
@@
$=\exp(-\rho_{1}(k)+\mathcal{O}(1))\mathbb{P}\!\left(S_{k}>t\right)$
$\geq\exp(-\rho_{1}(\varphi_{1}^{-1}(t))+\mathcal{O}(1))\mathbb{P}\!\left(S_{k}-\varphi_{1}(k)>0\right)$
$=\exp(-d^{*}t-\mathcal{K}_{\alpha}(t)+\mathcal{O}(1))(\mathbb{P}\!\left(M_{\infty}>0\right)+o(1)),$
@@
$M=M(t)\coloneq\lceil(1-\varepsilon)t\rceil\qquad\text{and}\qquad k_{i}\coloneq\lceil\varphi_{1}^{-1}(t-i)\rceil\quad\text{for }i\in\{0,\ldots,M\}.$ (5.17)
@@
As $\theta(1-\varepsilon)>d^{*}$ by the choice of $\theta$ and $\varepsilon$, and $\mathcal{K}_{\alpha}=o(t)$ by Lemma 5.5(c)(c), it follows that the second term is negligible compared to the first term. To deal with the sum, we define
$\widetilde{\alpha}(k)=\sum_{i=0}^{k-1}\frac{|d(i)-d^{*}|}{b(i)+d(i)}\qquad\text{for }k\in\mathbb{N}_{0},\quad\text{and}\quad\mathcal{K}_{\widetilde{\alpha}}(t)\coloneq\widetilde{\alpha}\big(\varphi_{1}^{-1}(t)\big)\qquad\text{for }t\geq 0,$ (5.21)
@@
We note that this is possible since $\mathcal{K}_{\widetilde{\alpha}}(t)=o(t)$ (this follows in the same way as $\mathcal{K}_{\alpha}(t)=o(t)$ in Lemma 5.5(c)(c)) and $M\geq(1-\varepsilon)t$. Using the second part of Lemma 5.5(c)(c) yields
@@
$\mathcal{K}_{\alpha}(t-(i+1))\leq\sup_{x\in[\varepsilon t,t]}|\mathcal{K}_{\alpha}(x)|\leq\sup_{x\in[\varepsilon t,t]}\mathcal{K}_{\widetilde{\alpha}}(x)=\mathcal{K}_{\widetilde{\alpha}}(t),$ (5.26)
@@ 5.2.2. Non-survival of old individuals
$\ell_{t}\coloneq\frac{d^{*}}{\lambda^{*}+d^{*}}t+\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))-K,\qquad u_{t}\coloneq\frac{d^{*}}{\lambda^{*}+d^{*}}t+\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))+K.$ (5.30)
@@
$\exp\big(\lambda^{*}{}$ $\ell_{t}-d^{*}(t-A-\ell_{t})-\mathcal{K}_{\alpha}(t-A-\ell_{t})+C_{1}\big)$ (5.33)
$=\exp\Big({}$ $\mathcal{K}_{\alpha}(r(t))-\mathcal{K}_{\alpha}\Big(\frac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))+K-A\Big)-(\lambda^{*}+d^{*})K+d^{*}A+C_{1}\Big),$
where $C_{1}>0$ is a sufficiently large constant. We use Lemma 5.5(c)(c) to obtain that for any $K,A>0$ there exists $t^{\prime}=t^{\prime}(K,A)>0$ such that for all $t\geq t^{\prime}$,
@@
Then, Assumption $\mathcal{K}_{\alpha}$ yields that there exists a constant $C_{2}>0$ such that for all $t>0$,
@@
$\mathbb{P}_{\mathcal{S}}\left(O^{\mathrm{ct}}_{t-A}\leq\ell_{t}\right)\leq\exp\big(C_{2}+1-(\lambda^{*}+d^{*})K+d^{*}A+C_{1}\big)+\varepsilon/4\leq\varepsilon/2,$ (5.36)
where we choose $K$ sufficiently large so that the exponential term is at most $\varepsilon/4$ in the final step. We note that we can first choose such a $K$ and then take $t^{\prime}=t^{\prime}(K,A)$ sufficiently large such that (5.34) holds.
@@
$\mathcal{E}_{t}\coloneq\{\forall v\in\mathcal{B}(\widetilde{u}_{t},u_{t}):L^{(v)}\leq t+A-\widetilde{u}_{t}\},\qquad\mathcal{D}_{t}\coloneq\{|\mathcal{B}(\widetilde{u}_{t},u_{t})|>m\}\quad\text{for }m\in\mathbb{N},t\geq 0.$ (5.38)
@@
$\mathbbm{1}_{\mathcal{D}_{t}}\mathbb{P}\!\left(\mathcal{E}_{t}\,|\,\mathcal{F}_{\widetilde{u}_{t}},|\mathcal{B}(\widetilde{u}_{t},u_{t})|\right)\leq\mathbbm{1}_{\mathcal{D}_{t}}\big(1-\mathbb{P}\!\left(L\geq t+A-\widetilde{u}_{t}\right)\big)^{m}\leq\exp\big(-m\mathbb{P}\!\left(L\geq t+A-\widetilde{u}_{t}\right)\big),$ (5.41)
@@
We choose $m=M^{-1}\exp(\lambda^{*}u_{t})$ for a large constant $M>0$ to be determined later to obtain the upper bound
@@
where $C_{3}>0$ is a large constant. We again use Lemma 5.5(c)(c) and Assumption $\mathcal{K}_{\alpha}$ to obtain similar bounds as in (5.34) and (5.35). We thus obtain, for some constants $C_{4}>0$ and $t^{\prime}=t^{\prime}(K,A)>0$ that for all $t>t^{\prime}$ we have have the upper bound
$\mathbb{P}(\mathcal{S})^{-1}\exp\Big(-M^{-1}\exp\Big(C_{4}+(d^{*}/2+\lambda^{*})K-d^{*}A-C_{3}\Big)\Big).$ (5.45)
Given $A$, we bound $C_{4}-d^{*}A-C_{3}\geq-C_{5}$ for some large constant $C_{5}=C_{5}(A,d^{*})$¿0. Then, we choose $K$ sufficiently large so that the entire term is at most $\varepsilon/4$. Again, we can first choose such a $K$ and then choose $t^{\prime}$.
@@
$\leq{}$ $\mathbb{P}_{\mathcal{S}}\left(|\mathcal{B}(0,u_{t})|\mathrm{e}^{-\lambda^{*}u_{t}}\leq\frac{2}{M}\right)$
$+\mathbb{P}_{\mathcal{S}}\left(|\mathcal{B}(0,\widetilde{u}_{t})|\mathrm{e}^{-\lambda^{*}u_{t}}\geq\frac{1}{M}\mathrm{e}^{\lambda^{*}\frac{K}{2}}\right).$
@@ 5.3. The richest individual
In this sub-section we provide an asymptotic expansion of $I^{\mathrm{ct}}_{t}$, the birth-time of the richest individual in $\mathrm{BP}(t)$. That is, the individual that is alive at time $t$ and has produced the largest offspring by time $t$. To this end, we show that individuals that are ‘too young’ produce ‘small’ offspring, whereas there exist ‘old’ individuals that produce large offspring. Combined, these then imply that $I_{t}^{\mathrm{ct}}$ is close to $O_{t}^{\mathrm{ct}}$.
@@ 5.3.1. Young individuals have small offspring and no individual has too large offspring
We start by showing that individuals that are alive in an interval of time centred at $t$ but born ‘too late’, that is, born too long after the birth-time $O_{t}^{\mathrm{ct}}$ of the oldest alive individual (at time $t$), have too small offspring to attain the largest offspring among all alive individuals. Similarly, we show that there does not exist any individual that attains a ‘too large’ offspring by time $\approx t$.
@@
$k$ $\coloneq\lceil\varphi_{1}^{-1}\big(\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\tfrac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))-C_{1}\big)\rceil,\qquad u_{t}\coloneq\frac{d^{*}}{\lambda^{*}+d^{*}}t+\frac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t)),$ (5.50)
$k^{\prime}$ $\coloneq\lceil\varphi_{1}^{-1}\big(\tfrac{\lambda^{*}}{\lambda^{*}+d^{*}}t-\tfrac{1}{\lambda^{*}+d^{*}}\mathcal{K}_{\alpha}(r(t))+C_{1}^{\prime}\big)\rceil,$
where we recall the function $r$ from Assumption $\mathcal{K}_{\alpha}$. Then, we have the following result.
@@
$\mathbb{P}_{\mathcal{S}}\left(\forall s\in[t-A,t+A]\ \forall v\in\mathcal{B}(u_{t}+C_{2},t)\cap\mathcal{A}_{s}^{\mathrm{ct}}:\deg_{s}(v)<k\right)\geq 1-\varepsilon.$ (5.51)
@@
$\mathbb{P}_{\mathcal{S}}\left(\forall s\in[t-A,t+A]\ \forall v\in\mathcal{A}_{s}^{\mathrm{ct}}:\deg_{s}(v)<k^{\prime}\right)\geq 1-\varepsilon.$ (5.52)
The first result states that all individuals born ‘too late’, namely after time $u_{t}+C_{2}$, that survive up to time $s\in[t-A,t+A]$, have a degree smaller than $k$ with probability close to $1$. The second result tells us that no individual that survives up to time $s\in[t-A,t+A]$ has degree larger than $k^{\prime}$ with probability close to one.
@@
$\mathbb{P}_{\mathcal{S}}{}$ $\big(\exists s\in[t-A,t+A]\ \exists v\in\mathcal{B}(u_{t}+C_{2},t)\cap\mathcal{A}_{s}^{\mathrm{ct}}:\deg_{s}(v)\geq k\big)$ (5.53)
@@
We now introduce the event $\{\sup_{t\geq 0}|\mathcal{B}(0,t)|\mathrm{e}^{-\lambda^{*}t}\leq M\}$, for some large but fixed $M>0$ to be determined. We then partition the set $\mathcal{B}(u_{t},t)$ into sets of individuals born in the intervals $(u_{t}+C_{2}+(j-1),u_{t}+C_{2}+j)$, i.e. into sets $\mathcal{B}(u_{t}+C_{2}+(j-1),u_{t}+C_{2}+j)$, with $j\in[\lceil t-u_{t}-C_{2}\rceil]$ and apply a union bound. For individuals $v\in\mathcal{B}(u_{t}+C_{2}+(j-1),u_{t}+C_{2}+j)$, we can bound their birth-time $\mathcal{B}(v)$ from below and above by $u_{t}+C_{2}+(j-1)$ and $u_{t}+C_{2}+j$, respectively. Similarly we can bound $s$ from below and above by tAt-A and $t+A$, respectively. As a result, we bound the above probability from above by
@@
We now choose $M$ large enough so that, by (5.13) in Corollary 5.4, the probability on the second line is at most $\varepsilon/2$. It thus remains to bound the sum from above by $\varepsilon/2$ to arrive at the desired result. For any $j\in[\lceil t-u_{t}\rceil]$, we omit the part of the event that $S_{D+1}>t-A-(u_{t}+C_{2}+j)$, use the independence of $D$ and $S_{k}$ and use a Chernoff bound with $\theta>0$ to arrive at the upper bound
@@
$\log\mathbb{E}\left[\mathrm{e}^{-\theta S_{k}}\right]=\sum_{i=0}^{k-1}\log\Big(1+\frac{\theta}{b(i)+d(i)}\Big)\geq\theta\varphi_{1}(k)-\frac{\theta^{2}}{2}\varphi_{2}(k).$ (5.56)
Since $\varphi_{2}(k)$ is increasing and converges to $\varphi_{2}(\infty)\coloneq\lim_{k\to\infty}\varphi_{2}(k)$ as $k$ tends to infinity by Assumption (C-V), and by using Lemma 5.1, we finally arrive at the upper bound
@@
We apply Lemma 5.5(c)(c) and Assumption $\mathcal{K}_{\alpha}$ and set $\theta=(C_{2}+(j-1)-(A+C_{1}))/\varphi_{2}(\infty)$ (which is positive for all $j\geq 1$ when $C_{2}$ is large), to bound this for all $t>t^{\prime}$ by
@@
where $C>0$ and $t^{\prime}$ are fixed constants, where $C$ is independent of all other constants and parameters, and $t^{\prime}=t^{\prime}(C_{1},d^{*},\lambda^{*})$. By substituting this upper bound in the sum in (5.54), we can bound it from above by